OpenStudy (liliakarina):

Determine the coordinates of thw midpoint between: K(-3,-4) and L(-5,-1)

OpenStudy (amistre64):

mid point is mid way is the average of the 2 points is (K+L)/2

OpenStudy (liliakarina):

I see

OpenStudy (amistre64):

i hope thats not sarcasm cause i cant read that to well on this thing lol

OpenStudy (amistre64):

what K+L ?

OpenStudy (liliakarina):

It's okay, only that I have to use a specific formula

OpenStudy (amistre64):

this is the formula ...

OpenStudy (liliakarina):

Okay, thank you

OpenStudy (amistre64):

add up like parts and divide in half K(-3,-4) +L(-5,-1) --------- (-8,-5)/2 = (-4, -5/2)

OpenStudy (amistre64):

\[\left(\frac{Kx+Lx}{2},\frac{Ky+Ly}{2}\right)\]

OpenStudy (liliakarina):

Excellent, thanks a lot!

OpenStudy (amistre64):

yw